how many grams of carbon are neede to completely react with 125 g of iron3 oxide?
1
Expert's answer
2013-03-29T04:14:56-0400
The chemical equation for this reaction is

2Fe2O3 + 3C = 4Fe + 3CO2

n(mol) = m(g) / MW(g/mol)

n(Fe2O3) = 125 / (55.8 x 2 + 16 x 3) = 0.783 mol

According to the chemical equation n(C) = n(Fe2O3) : 2 x 3 = 0.783 : 2 x 3 =1.17 mol

m(C) = n x MW = 1.17 x 12 = 14.04 g

Answer: m(C) = 14.04 g